OpenStudy (anonymous):

AGAIN HELP ME!! lim cos (x+1/2 pi)/ x approaches x-> 0

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

do you get to use l'hopital's rule, or something else?

OpenStudy (misty1212):

if you cannot use it, then recognize this as the derivative of \[\cos(x+\frac{\pi}{2})\] evaluated at \(x=0\)

OpenStudy (misty1212):

probably easier to do it the second way in any case do you know the derivative of cosine?

OpenStudy (anonymous):

Actually Im kind of confused ! S: could show me step by step

OpenStudy (misty1212):

lets see what you are allowed to use first do you know derivatives?

OpenStudy (amistre64):

i was considering power series ... :)

OpenStudy (misty1212):

lol

OpenStudy (misty1212):

i have a better idea, lets use a trig identity, then it will take no steps ok one step

OpenStudy (misty1212):

\[\cos(x+\frac{\pi}{2})=-\sin(x)\] so you have \[-\lim_{x\to 0}\frac{\sin(x)}{x}\]

OpenStudy (misty1212):

that limit is supposed to be well known

OpenStudy (anonymous):

Oh thanks!!!! I got it :DDDD

OpenStudy (misty1212):

did you get \(-1\) ?

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

:D yes